OpenStudy (anonymous):

The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 1120. If three times the number of one-step equations minus twice the number of two-step equations is equal to 1300, how many two-step equations eliminated today? I got 412

OpenStudy (anonymous):

o = one step equations t = two step equations o + t = 1120 -->o = 1120 - t 3o - 2t = 1300 3(1120 - t) - 2t = 1300 3360 - 3t - 2t = 1300 -5t = 1300 - 3360 -5t = -2060 t = 412 yep...thats what I got

OpenStudy (anonymous):

x=412 y=708

OpenStudy (anonymous):

so 412 thanks guys

OpenStudy (anonymous):

can you help me with one more question @tcarroll010

OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + 2y = 12 x - y = 3

OpenStudy (anonymous):

do you prefer substitution or elimination ?

OpenStudy (anonymous):

substitution

OpenStudy (anonymous):

wiat is it 2

OpenStudy (anonymous):

x - y = 3 ---> x = y + 3 now just sub y + 3 in for x in the 2nd equation you can do this...you did the first one. Tell me what you get and I will do it and see if we get the same answer.

OpenStudy (anonymous):

x is 2

OpenStudy (anonymous):

@kelliegirl33 ???

OpenStudy (anonymous):

sorry....I am back...let me check...hold on

OpenStudy (anonymous):

4x + 2y = 12 4(y + 3) + 2y = 12 4y + 12 + 2y = 12 6y = 0 y = 0 x = y + 3 x = 0 + 3 x = 3 let me check it... 4x + 2y = 12 4(3) + 2(0) = 12 12 + 0 = 12 12 = 12 (correct) No...I got x = 3 and y = 0

OpenStudy (anonymous):

ok i see thanks a ton

OpenStudy (anonymous):

no problem....glad to help :)

OpenStudy (anonymous):

when in doubt...check it